Linda Jones 250 miles using 9 gallons of gas at this rate how many gallons of gas will she need to drive 275 miles
1
Expert's answer
2020-11-25T18:36:03-0500

9 gallons is used for 250 miles

Therefore, 1 gallon will be used for x miles

9 gallons = 250 miles

1 gallon = x miles

Crossmultiply

9x = 250

Therefore, 1 gallon will be used for (250/9) miles


1 gallon = (250/9) miles

y gallons = 275 miles

Crossmultiply

(250/9)y = 275

y = 275/(250/9)

y = 9.9 gallons.

Therefore, 9.9 gallons will be needed to drive 275 miles.